Business School Problem 3 (25 points) Shamrock Inc. are Shamrocks purchases during Juy sells high-definition televisions and uses a perpetual inventory system. Shown below Purchase Date Unit purchased July 3 July 7 Unit cost 190 195 Total cost $1,710 $2,340 Total 32 $6,305 On July 20, Shamrock Inc. sold 15 units. The other 17 units remained in inventory at July 31 1. Compute the cost of goods sold relating to the sale on July 20 and the ending inventory July 31 using the following cost flow assumptions. (15 points) (Show your solutions, includling the number of units and costs per unit) Ending Cost of Inventory Goods Sold a Average cost b FIFO (First-in, first-out) с LIFO (Lastin, first-out) 2. Ulsing the cost figures computed above, answer the following questions: will result in the company reporting the lowest net a. Which of the three cost flow assumptions income for the current year? Would this always be the case? Explain. (5 points b. Which of the three cost flow assumptions will result in the highest income tax expense for the year?Would you expect this usualy to be the case? Explain. (5 points) www.euruni.edu
Answer #1
c) COGS Ending Inventory
a) Average cost:
Average unit cost = (1710+2340)/21 = $192.86
COGS = 15*192.86 = $      2,892.90
= 6305-2892.90 = $ 3,412.10
b) FIFO:
COGS = 9*190+6*195 = $      2,880.00
= 6305-2880 = $ 3,425.00
c) LIFO:
OGS = 12*195+3*190 = $      2,910.00
= 6305-2910 = $ 3,395.00
2) The assumption with highest COGS will report the lowest NI = LIFO.
The need not always be the case. It will not hold good if the unit prices are
on the decrease for successive purchases.
3) The assumption with the lowest COGS will report the highest NI and highest
tax. It is FIFO. It need not always be the case. In cases of decreasing prices,
FIFO will show highest COGS and lowest NI and lowest tax expense.